OpenStudy (anonymous):

Solve. (x + 4)^2 = 4

OpenStudy (anonymous):

x + 4 = |2| x = -2 - 4 = -6 or x = 2 - 4 = -2

OpenStudy (anonymous):

would the other answer be -6?

OpenStudy (anonymous):

ohhh! haha i see! thank you.<3

OpenStudy (anonymous):

he skipped the quadratic equation part

OpenStudy (anonymous):

-6 and -2

OpenStudy (anonymous):

Oh -skipper-

OpenStudy (cornzyblack):

Naaa there are two solns. U either use his(Ishaan) method or Quadratic eqn

OpenStudy (anonymous):

yeah he just square rooted both sides which works out on this luckily.